A pre-filtered river water 10ml is added to a 100ml volumetric flask along with the 20ml of La+cs solution (to minimise matrix effects) and then ultrapure water is added to make up the mark.


A Ca²+ concentration of 2.3 meq/L is determined from the calibration curve . 


What is the Ca²+ concentration in the original river water sample in meq/L and mg/L?


1
Expert's answer
2021-04-21T03:52:33-0400

Molarity = mole of the solute/total Solution

Molar Mass of Ca = 12.01

= 20/(10+100 + 20)

= 20/130×2.3

= 0.3538 × 12

= 4.2462m/L